# 3D Unsupervised Learning by Distilling 2D Open-Vocabulary Segmentation Models for Autonomous Driving

### v1

@@ Abstract
Point cloud data labeling is considered a time-consuming and expensive task in autonomous driving, whereas unsupervised learning can avoid it by learning point cloud representations from unannotated data. In this paper, we propose UOV, a novel 3D Unsupervised framework assisted by 2D Open-Vocabulary segmentation models. It consists of two stages: In the first stage, we innovatively integrate high-quality textual and image features of 2D open-vocabulary models and propose the Tri-Modal contrastive Pre-training (TMP). In the second stage, spatial mapping between point clouds and images is utilized to generate pseudo-labels, enabling cross-modal knowledge distillation. Besides, we introduce the Approximate Flat Interaction (AFI) to address the noise during alignment and label confusion. To validate the superiority of UOV, extensive experiments are conducted on multiple related datasets. We achieved a record-breaking 47.73% mIoU on the annotation-free point cloud segmentation task in nuScenes, surpassing the previous best model by 10.70% mIoU. Meanwhile, the performance of fine-tuning with 1% data on nuScenes and SemanticKITTI reached a remarkable 51.75% mIoU and 48.14% mIoU, outperforming all previous pre-trained models. Our code is available at: https://github.com/sbysbysbys/UOV.
1 Introduction
@@
As neural-network-based 3D scene perception methods, e.g., object detection [1, 2, 3, 4], point cloud segmentation [5, 6, 7], etc., become increasingly complex in their network architectures with a growing number of parameters, methods relying solely on enhancing model structures are reaching a point of saturation. Meanwhile, approaches to enhancing model performance through data-driven methods heavily rely on time-consuming and expensive manual annotations. Due to constraints such as insufficient class annotations, applying traditional point cloud perception methods to large-scale unlabeled data meets significant challenges.
@@
(a) Segmentation results of UOV annotation-free training. More illustrations are presented in Fig. 5
(b) Results of various tasks on nuScenes
Figure 1: Performance of UOV on nuScenes [8]. Our annotation-free segmentation approach demonstrates comparable results to some supervised methods (Fig. 1a), while achieving optimal performance across various tasks on nuScenes (Fig. 1b).
@@
Unsupervised learning is a powerful machine learning paradigm that enables representation learning from data without the need for manually annotated supervision signals. Multi-modal unsupervised algorithms for 3D scene perception tasks integrate the internal structure of point clouds with image or text knowledge to generate objectives. They bypass the costly manual annotations, bridging the gap between traditional 3D perceptual models and unannotated data.
@@
Existing 3D unsupervised methods for annotation-free training [9, 10, 11, 12] aim to transfer knowledge from visual foundation models, e.g., Contrastive Vision-Language Pre-training (CLIP) [13] or Segment Anything (SAM) [14], to point cloud representations. However, 3D unsupervised models based on CLIP [13] suffer from intolerable noise, while SAM [14] fails to correspond texts and images. Therefore, we seek high-quality image segmentation models with textual correspondences to serve as teacher models for 3D unsupervised learning. Recently, CLIP-based 2D open-vocabulary segmentation models [15, 16, 17, 18] have demonstrated excellent performances. They employ contrastive learning to extract textual and image features from a shared embedding space and are capable of segmenting and identifying objects from a set of open classes in various environments. These models provide us with image segmentation and labels corresponding to the segmented regions, as well as easy-to-extract text and image representations. Meanwhile, they significantly outperform other Visual Language Models (VLMs) [13, 19, 20, 21] in open-vocabulary segmentation tasks, as shown in Fig. 3, 6.
@@
In this paper, we propose UOV, a 3D Unsupervised framework by distilling 2D Open-Vocabulary segmentation models. It aims to address the difficulties of point cloud perception using unannotated data through 3D unsupervised learning. Previous work [12, 9] utilizes contrastive learning between point clouds and texts for distilling knowledge, which improves the model’s scene understanding ability but the performance cannot be guaranteed. Hence, one can conclude that using only contrastive learning as the method for annotation-free training is insufficient. Based on this premise, UOV adopts a novel two-stage unsupervised strategy. The first stage uses Tri-Modal contrastive Pre-training (TMP) to warm up the network parameters, where we innovatively incorporate the text modality to enhance the semantic perception of point cloud representations. The training objectives in contrastive loss are all derived from 2D open-vocabulary segmentation models and synchronous generated with segmentations and labels. The second stage is pseudo-label guided annotation-free training, which maps the image segmentation from the 2D open-vocabulary segmentation model to point cloud pseudo-labels, completing the knowledge distillation from 2D to 3D. Meanwhile, it refines semantic comprehension by using a class dictionary, which establishes mappings from open vocabularies to semantic classes. Furthermore, to address the perceptual limitations of images in 3D scenes, such as noise during alignment and label confusion, we introduce Approximate Flat Interaction (AFI), which corrects the error through point cloud spatial interaction over a wide perceptual domain.
@@
In the experiment, We selected four CLIP-based open-vocabulary segmentation models, ie., MaskCLIP [15], FC-CLIP [16], SAN [17], CAT-Seg [18], as teacher models. To validate the performance of our method in unsupervised point cloud segmentation tasks, extensive experiments on multiple autonomous driving datasets were conducted. Firstly, UOV achieved a remarkable improvement of 10.70% mIoU, reaching a Top-1 accuracy of 47.74% mIoU in benchmark testing of annotation-free segmentation performance on nuScenes [8]. Furthermore, to compare the advantages of our approach with previous pre-training methods, we conducted 1% data fine-tuning and 100% data linear-probing experiments on nuScenes, yielding mIoU scores of 51.75% and 56.35%. In comparison to the current best pretraining method, our approach achieved improvements of 5.91% and 11.95% mIoU, respectively. When fine-tuning with 1% data on SemanticKITTI, UOV achieved a 48.14% mIoU, surpassing the previous best method by 1.51% mIoU. UOV demonstrated state-of-the-art performance across various experiments, validating its effectiveness.
Overall, the key contributions of our work are summarized as follows:
1. To the best of our knowledge, our approach not only represents the first attempt to comprehensively utilize state-of-the-art 2D open-vocabulary segmentation models for training 3D unsupervised networks but also the first to synchronously generate knowledge before pre-training and incorporate text featues into pre-training.
2. We propose a two-stage unsupervised framework, UOV, which leverages 2D open-vocabulary segmentation models to assist in unsupervised training without annotation. Additionally, a detachable spatial interaction method, AFI, is introduced to enhance the spatial understanding of the model.
@@
3. Our approach not only breaks through in annotation-free semantic segmentation (Fig. 1a), but also notably outperforms prior state-of-the-art methods in other downstream tasks (Fig. 1b).
2 Related Work
2.1 CLIP-based 2D Open-Vocabulary Segmentation
@@
2D open-vocabulary segmentation models aim to segment all categories in the real world. Traditional open-vocabulary image segmentation models [22, 23, 24] attempt to learn image embeddings aligned with text embeddings. Inspired by Visual Language Models (VLMs), e.g., CLIP [13] and ALIGN [25], which have demonstrated remarkable performance in 2D tasks, recent studies have attempted to transfer CLIP’s outstanding zero-shot segmentation capability to open-vocabulary tasks [17, 18, 15, 16, 26, 27, 28, 29, 30]. In notable works, LSeg [26] learns pixel-level visual embeddings from CLIP, marking the first exploration of CLIP’s role in language-driven segmentation tasks. More recently, MaskCLIP [15] obtains pixel-level embeddings by modifying the CLIP image encoder. SAN [17] augments CLIP with lightweight side networks to predict mask proposals and categories. CAT-Seg [18] proposes a cost-aggregation-based method to optimize the image-text similarity map. Additionally, FC-CLIP [16] integrates all components into a single-stage framework using a shared frozen convolutional CLIP backbone. These works utilize CLIP as the central component of their network, granting them robust segmentation and recognition capabilities, as well as an image-text-aligned structure. Consequently, they can provide us with high-quality knowledge.
2.2 3D Unsupervised Learning in Large-scale Point Clouds
@@
Unsupervised learning can be utilized for learning point cloud representations. Mainstream 3D unsupervised pre-training methods are mainly reconstruction-based [31, 32, 33, 34], or contrast-based [35, 36, 37, 38, 39, 40, 41]. However, many of these methods are constrained by the quantity of point clouds, limiting their applicability to single-object or indoor scene learning. Prior attempts, such as PointContrast [36], DepthContrast [37], SegContrast [42], and PPKT [38], have built contrastive objectives on large-scale point clouds. Additionally, SLidR [39] adopts a novel approach by leveraging a superpixel-superpoint correspondence for 3D-to-2D spatial alignment, showing promising performance on autonomous driving datasets. Built upon SLidR, SEAL [41] employs VLMs to aid in superpixel generation.
@@
Recently, inspired by the achievements of CLIP [13], numerous works have focused on reproducing the excellent performance demonstrated by CLIP in 3D, not limited to pre-training. In 3D scene understanding, CLIP2Scene [9], similar to OpenScene [12], embeds the knowledge of CLIP feature space into representations of 3D point cloud, enabling annotation-free point cloud segmentation. PLA [43] and RegionPLC [44] accomplish scene understanding through point-language alignment or contrastive learning framework. CLIP2 [10] demonstrates perfect zero-shot instance segmentation performance through language-3D alignment at the semantic level and image-3D alignment at the instance level. Unlike the others, Chen et al. [11] utilizes CLIP to generate pseudo-labels and uses SAM [14] to assist in denoising.
In general, most of these methods rely on either feature alignment or pseudo-labeling. However, using only contrastive learning may incorrectly pull apart positive samples of the same class when employing point-level correspondence methods, e.g., pixel-point, pixel to FPS and $k$-NN. It is also susceptible to differences in data volume for each class and asymmetric point cloud densities. Meanwhile, the use of pseudo-labeling is greatly limited by the accuracy of the teacher’s model. To address these issues, our work employs state-of-the-art 2D open-vocabulary segmentation models as teacher models. We not only distill knowledge onto point cloud using pseudo-labels, but also incorporate multi-modal contrastive learning into our framework.
3 Method
@@
3.1 Extracting Knowledge from CLIP-based 2D Open-Vocabulary Segmentation Models
@@
In perceptual approaches to unknown classes in 2D, unlike zero-shot learning, open-vocabulary learning uses language data as supervision. In terms of network structure, MaskCLIP [15] changes the image encoder of CLIP to propose pixel-level representations instead of image-level. SAN [17] proposes a side adapter network attached to a frozen CLIP encoder; CAT-Seg [18] employs a cost-aggregation-based method to improve CLIP; FC-CLIP [16] adds a decoder, mask generator, in-vocabulary classifier, and out-vocabulary classifier after freezing the CLIP backbone. In most cases, the mask generator operates independently of the class generator. Pixel-level features generated by the modified CLIP-based network are max-pooled for each mask, and the objective loss is computed with the text features.
@@
The aforementioned methods hold a similar view with contrastive learning for point cloud pre-training. The difference is that the latter uses image-point cloud contrastive learning [38, 39] (some of the work uses data augmentation for single-modal contrastive learning). Most of them use SLIC [45], SAM [14], and SEEM [19] to guide mask segmentation and choose ResNet [46] as the image encoder, which means that mask segmentation and mask feature generation are two completely independent modules. This not only increases the training time but also makes noise easily stack up across different models. At the same time, the lack of language guidance during segmentation will lead to a more random mask granularity. Fortunately, 2D open-vocabulary segmentation models perfectly address this issue, as we can not only extract labels and image embeddings from them but also obtain segmentation with appropriate granularity.
@@
Figure 2: Overview of UOV, which consists of two stages: Tri-Modal Pre-training (TMP) and Annotation-free training (UOV-baseline). Both stages leverage masks $M_{\mathcal{I}}$ and mask labels $L_{M}$ extracted from 2D open-vocabulary segmentation models, while mask features $F_{M}$ and text features $F_{T}$ are employed only in TMP. TMP enhances scene understanding through contrastive losses: superpixel-superpoint loss $\mathcal{L}_{I-P}$ and text-superpoint loss $\mathcal{L}_{T-P}$, while our baseline employs pseudo-labels to supervise the 3D network. Additionally, to bridge dataset classes and open vocabularies, we introduce a class dictionary $\mathcal{C}$. The Approximate Flat Interaction (AFI) optimizes the results of annotation-free training by spatial structural analysis in a broad perception domain.
@@
To summarize, we extract four interrelated, synchronously generated knowledge from CLIP-based open-vocabulary segmentation models: 1) images’ segmentations as masks $M_{\mathcal{I}}$ from image set $\mathcal{I}$; 2) corresponding labels $L_{M}$ for $M_{\mathcal{I}}$; 3) image features corresponding to $M_{\mathcal{I}}$, denoted as $F_{M}$; and 4) text features $F_{T}$. Each of the above knowledge will play an important role in Sec. 3.2,3.3,3.4, as shown in Fig. 2.
3.2 Baseline of UOV
@@
Given a point cloud $\mathcal{P}=\{(p_{n},e_{n})|n=1,\ldots,N\}$, where $p_{n}\in\mathbb{R}^{3}$ represents the 3D coordinates of a point, $e_{n}\in\mathbb{R}^{L}$ denotes the point’s features. ${L}=\{l_{n}|n=1,\ldots,N\}$ are the labels of $\mathcal{P}$ and =$\mathcal{I}=\{i_{k}|k=1,\ldots,K\}$ represents the images captured by a synchronized camera at the same moment. In contrast to supervised methods, our task does not utilize labels ${L}$ during training. We choose to employ a simple way of generating pseudo-labels for point clouds with the assistance of image segmentation: With masks $M_{\mathcal{I}}=\{m_{r}|r=1,\ldots,R\}$ obtained from image set $\mathcal{I}$ as described in Sec. 3.1, we use the labels ${L}_{M}$ corresponding to $M_{\mathcal{I}}$ as the pseudo-label ${L}_{\text{pixel}}^{\text{pseudo}}$ for pixels in every mask $m_{r}\in M_{\mathcal{I}}$. By leveraging known sensor calibration parameters, we establish a mapping $\Gamma_{\text{camera}\leftarrow\text{LiDAR}}$ to bridge the gap between domains of point clouds and images. Pseudo-labels ${L}_{\mathcal{P}}^{\text{pseudo}}=\{l_{n_{0}}^{\text{pseudo}}|n_{0}=1,\ldots,N_{0}\}$ for point clouds $\mathcal{P}$ are generated through ${L}_{\text{pixel}}^{\text{pseudo}}$ and mapping $\Gamma_{\text{camera}\leftarrow\text{LiDAR}}$. For a 3D backbone $\mathcal{F}_{\theta_{p}}:\mathbb{R}^{N\times(3+L)}\rightarrow\mathbb{R}^{N\times D}$ with the learnable parameter $\theta_{p}$, we train $\theta_{p}$ with pseudo-labels ${L}_{\mathcal{P}}^{\text{pseudo}}$. Given the sparsity of point clouds, it is obvious that $\Gamma_{\text{camera}\leftarrow\text{LiDAR}}$ is not surjective. It is important to note that $\Gamma_{\text{camera}\leftarrow\text{LiDAR}}$ is also not injective, as the projection area of LiDAR is not entirely covered by cameras, resulting in obvious pseudo-label-blank areas in point cloud $\mathcal{P}$. After knowledge distillation, these untrained regions exhibit label confusion, which will be discussed in Sec. 3.4.
@@
3.3 Tri-Modal Contrastive Pre-training (TMP)
In CLIP-based 2D open-vocabulary segmentation models, masks and image features mostly share encoders. They are generated synchronously and guided by the same text. On the contrary, to the best of our knowledge, all existing 3D pre-training methods [38, 39, 41] using contrastive learning between image and point cloud generate masks and mask features asynchronously, which causes noise aggregation between different backbones. To break this situation, we introduce Tri-Modal contrastive Pre-training (TMP). TMP innovatively integrates text features into pre-training and removes the 2D backbone through pre-generation of the features, demonstrating excellent performance in both annotation-free training and fine-tuning. The schematic illustration of TMP can refer to Fig. 2.
3.3.1 Superpixel-Superpoint Generation
The usage of superpixels and superpoints for segmentation in SLidR [39] is inspiring, which overturns previous works in 2D-3D contrastive learning methods at the point level. The superpoint-superpixel approach partly alleviates the issue of "self-conflict", which means different point clouds of the same semantics form negative samples in contrastive learning. However, as analyzed in Sec. 3.1, the "self-conflict" problem caused by the mask granularity randomness of VLMs [45, 14] has not been properly addressed. This issue arises due to the lack of textual assistance, e.g., windows and doors of a car should not be treated as negative examples at the granularity of the entire car. On the contrary, 2D open-vocabulary segmentation models [16, 17, 18] perfectly fit the granularity by incorporating textual information defined in class dictionary $\mathcal{C}$. Also, as shown in Fig. 3, the accuracy of the superpixels is greatly influenced by the models. Based on these observations, it is reasonable to generate superpixels using CLIP-based 2D open-vocabulary segmentation models.

### v3

# 3D Annotation-Free Learning by Distilling 2D Open-Vocabulary Segmentation Models for Autonomous Driving

@@ Abstract
Point cloud data labeling is considered a time-consuming and expensive task in autonomous driving, whereas annotation-free learning training can avoid it by learning point cloud representations from unannotated data. In this paper, we propose AFOV, a novel 3D Annotation-Free framework assisted by 2D Open-Vocabulary segmentation models. It consists of two stages: In the first stage, we innovatively integrate high-quality textual and image features of 2D open-vocabulary models and propose the Tri-Modal contrastive Pre-training (TMP). In the second stage, spatial mapping between point clouds and images is utilized to generate pseudo-labels, enabling cross-modal knowledge distillation. Besides, we introduce the Approximate Flat Interaction (AFI) to address the noise during alignment and label confusion. To validate the superiority of AFOV, extensive experiments are conducted on multiple related datasets. We achieved a record-breaking 47.73% mIoU on the annotation-free 3D segmentation task in nuScenes, surpassing the previous best model by 3.13% mIoU. Meanwhile, the performance of fine-tuning with 1% data on nuScenes and SemanticKITTI reached a remarkable 51.75% mIoU and 48.14% mIoU, outperforming all previous pre-trained models. Our code is available at: https://github.com/sbysbysbys/AFOV
Introduction
@@
As neural-network-based 3D scene perception methods, e.g., object detection (Shi, Wang, and Li 2019; Shi et al. 2020; Zhou and Tuzel 2018; Lang et al. 2019), point cloud segmentation (Qi et al. 2017a, b; Choy, Gwak, and Savarese 2019), etc., become increasingly complex in their network architectures with a growing number of parameters, methods relying solely on enhancing model structures are reaching a point of saturation. Meanwhile, approaches to enhancing model performance through data-driven methods heavily rely on time-consuming and expensive manual annotations. Due to constraints such as insufficient class annotations, applying traditional point cloud perception methods to large-scale unlabeled data meets significant challenges.
@@
Figure 1: Segmentation results of AFOV annotation-free training. More illustrations are presented in Appendix D.
Figure 2: Performance of AFOV on nuScenes.
@@
Annotation-free learning is a powerful machine learning paradigm that enables representation learning from data without the need for manual supervision. Multi-modal annotation-free algorithms for 3D scene perception tasks integrate the internal structure of point clouds with image or text knowledge to generate objectives. They bypass the costly manual annotations, bridging the gap between traditional 3D perceptual models and unannotated data.
@@
Existing 3D annotation-free methods (Chen et al. 2023b; Zeng et al. 2023; Chen et al. 2023a; Peng et al. 2023) aim to transfer knowledge from visual foundation models (VFMs), e.g., Contrastive Vision-Language Pre-training (CLIP) (Radford et al. 2021) or Segment Anything (SAM) (Kirillov et al. 2023), to point cloud representations. However, 3D annotation-free models based on CLIP (Radford et al. 2021) suffer from intolerable noise, while SAM (Kirillov et al. 2023) fails to correspond texts and images. Therefore, we seek high-quality image segmentation models with textual correspondences to serve as teacher models for 3D annotation-free learning. Recently, CLIP-based 2D open-vocabulary segmentation models (Zhou, Loy, and Dai 2022; Yu et al. 2023; Xu et al. 2023; Cho et al. 2023) have demonstrated excellent performances. They employ contrastive learning to extract textual and image features from a shared embedding space and are capable of segmenting and identifying objects from a set of open classes in various environments. These models provide us with image segmentation and labels corresponding to the segmented regions, as well as easy-to-extract text and image representations. Meanwhile, they significantly outperform other Visual Language Models (VLMs) (Radford et al. 2021; Zou et al. 2024, 2023; Li et al. 2023) in open-vocabulary segmentation tasks.
@@
In this paper, we propose AFOV, a 3D Annotation-Free framework by distilling 2D Open-Vocabulary segmentation models. It aims to address the difficulties of point cloud perception using unannotated data through 3D annotation-free learning. AFOV adopts a novel two-stage strategy: The first stage uses Tri-Modal contrastive Pre-training (TMP) to warm up the network parameters, where we innovatively incorporate the textual information to enhance the semantic perception of point cloud representations; The second stage is pseudo-label guided annotation-free training, completing the knowledge distillation from 2D to 3D. Furthermore, to address the perceptual limitations of AFOV, such as noise during alignment and label confusion, we introduce Approximate Flat Interaction (AFI). It provides a robust error correction mechanism for the above two stages through point cloud spatial interaction.
@@
In the experiment, We selected four CLIP-based open-vocabulary segmentation models, ie., MaskCLIP (Zhou, Loy, and Dai 2022), FC-CLIP (Yu et al. 2023), SAN (Xu et al. 2023), CAT-Seg (Cho et al. 2023), as teacher models. To validate the performance of our method in annotation-free point cloud segmentation tasks, extensive experiments on multiple autonomous driving datasets were conducted. Firstly, AFOV achieved a remarkable improvement of 3.13% mIoU, reaching a Top-1 accuracy of 47.74% mIoU in benchmark testing of annotation-free 3D segmentation on nuScenes (Caesar et al. 2020). Furthermore, treating AFOV as a pre-trained model, we conducted 1% data fine-tuning and 100% data linear-probing experiments on nuScenes, yielding mIoU scores of 51.75% and 56.35%. Compared to the current best pretraining method, AFOV achieved improvements of 4.16% and 4.81% mIoU, respectively. When fine-tuning with 1% data on SemanticKITTI (pre-training on nuScenes), AFOV achieved a 48.14% mIoU. AFOV demonstrated state-of-the-art performance across various experiments, validating its effectiveness.
Compared to scene understanding work based on point clouds (such as OpenScene), the introduction of text information allows AFOV to generate pseudo labels for knowledge distillation. As a result, AFOV directly predicts outcomes, whereas scene understanding models match features between point clouds and text. As with fully supervised closed-set point cloud segmentation models, the directly predicted output is far superior to the output obtained through feature matching.
Unlike most previous annotation-free 3D segmentation models, all the knowledge for training AFOV comes from state-of-the-art 2D open-vocabulary segmentation models. These open-vocabulary segmentation models (such as FC-CLIP, CAT-Seg, and SAN) perform far better than other VLMs like maskCLIP, SAM, and SEEM. They can also extract masks, labels, and features, avoiding noise accumulation between different backbones. Therefore, AFOV presents significant advantages over previous related works.
In conclusion, the main contributions of this work are:
  • We introduce a novel and efficient two-stage annotation-free training framework, AFOV, which comprehensively utilizes state-of-the-art 2D open-vocabulary segmentation models for knowledge distilling.
  • AFOV innovatively introduces TMP and AFI, addressing the issues in previous works. Moreover, we introduce the superpixel-superpoint into annotation-free 3D segmentation for the first time.
@@
  • Experimentally, our approach not only breaks through in annotation-free semantic segmentation (Fig. 1), but also notably outperforms prior state-of-the-art methods in other downstream tasks (Fig. 2).
Related Work
CLIP-based 2D Open-Vocabulary Segmentation
@@
2D open-vocabulary segmentation models aim to segment all categories in the real world. Traditional open-vocabulary image segmentation models (Zhao et al. 2017; Xian et al. 2019; Bucher et al. 2019) attempt to learn image embeddings aligned with text embeddings. Inspired by Visual Language Models (VLMs), e.g., CLIP (Radford et al. 2021) and ALIGN (Jia et al. 2021), which have demonstrated remarkable performance in 2D tasks, recent studies have attempted to transfer CLIP’s outstanding zero-shot segmentation capability to open-vocabulary tasks (Xu et al. 2023; Cho et al. 2023; Zhou, Loy, and Dai 2022; Yu et al. 2023; Li et al. 2022a; Ghiasi et al. 2022; Ding et al. 2022; Xu et al. 2022; Liang et al. 2023). In notable works, LSeg (Li et al. 2022a) learns pixel-level visual embeddings from CLIP, marking the first exploration of CLIP’s role in language-driven segmentation tasks. More recently, MaskCLIP (Zhou, Loy, and Dai 2022) obtains pixel-level embeddings by modifying the CLIP image encoder; SAN (Xu et al. 2023) augments CLIP with lightweight side networks to predict mask proposals and categories; CAT-Seg (Cho et al. 2023) proposes a cost-aggregation-based method to optimize the image-text similarity map. Additionally; FC-CLIP (Yu et al. 2023) integrates all components into a single-stage framework using a shared frozen convolutional CLIP backbone. These works utilize CLIP as the central component of their network, granting them robust segmentation and recognition capabilities, as well as an image-text-aligned structure. Consequently, they can provide us with high-quality knowledge.
3D Representation Learning Based on 2D-to-3D Knowledge distillation
@@
Unsupervised learning can be utilized for learning point cloud representations. Mainstream 3D unsupervised pre-training methods are mainly reconstruction-based (Boulch et al. 2023; Lin and Wang 2022; Hess et al. 2023; Min et al. 2023), or contrast-based (Li et al. 2022b; Xie et al. 2020; Zhang et al. 2021; Liu et al. 2021; Sautier et al. 2022; Mahmoud et al. 2023; Liu et al. 2023). However, many of these methods are constrained by the quantity of point clouds, limiting their applicability to single-object or indoor scene learning. Prior attempts, such as PointContrast (Xie et al. 2020), DepthContrast (Zhang et al. 2021), SegContrast (Nunes et al. 2022), and PPKT (Liu et al. 2021), have built contrastive objectives on large-scale point clouds. Additionally, SLidR (Sautier et al. 2022) adopts a novel approach by leveraging a superpixel-superpoint correspondence for 3D-to-2D spatial alignment, showing promising performance on autonomous driving datasets. Built upon SLidR, SEAL (Liu et al. 2023) employs VLMs to aid in superpixel generation.
@@
Recently, inspired by the achievements of CLIP (Radford et al. 2021), numerous works have focused on reproducing the excellent performance demonstrated by CLIP in 3D annotation-free tasks, not limited to pre-training. In 3D scene understanding, CLIP2Scene (Chen et al. 2023b), similar to OpenScene (Peng et al. 2023) and OV3D (Jiang, Shi, and Schiele 2024), embeds the knowledge of CLIP feature space into representations of 3D point cloud, enabling annotation-free point cloud segmentation; PLA (Ding et al. 2023) and RegionPLC (Yang et al. 2023) accomplish scene understanding through point-language alignment or contrastive learning framework; VLM2Scenes (Liao, Li, and Ye 2024) exploits the potential of VLMs; and CLIP2 (Zeng et al. 2023) demonstrates perfect zero-shot instance segmentation performance through language-3D alignment at the semantic level and image-3D alignment at the instance level. Unlike the others, Chen et al. (Chen et al. 2023a) utilizes CLIP to generate pseudo-labels and uses SAM (Kirillov et al. 2023) to assist in denoising.
Method
@@
Extracting Knowledge from CLIP-based 2D Open-Vocabulary Segmentation Models
@@
In perceptual approaches to unknown classes in 2D, unlike zero-shot learning, open-vocabulary learning uses language data as supervision. In terms of network structure, MaskCLIP changes the image encoder of CLIP to propose pixel-level representations instead of image-level. SAN proposes a side adapter network attached to a frozen CLIP encoder; CAT-Seg employs a cost-aggregation-based method to improve CLIP; FC-CLIP adds a decoder, mask generator, in-vocabulary classifier, and out-vocabulary classifier after freezing the CLIP backbone. In most cases, the mask generator operates independently of the class generator. Pixel-level features generated by the modified CLIP-based network are max-pooled for each mask, and the objective loss is computed with the text features.
@@
The aforementioned methods hold a similar view with contrastive learning for point cloud pre-training. The difference is that the latter uses image-point cloud contrastive learning (Liu et al. 2021; Sautier et al. 2022) (some of the work uses data augmentation for single-modal contrastive learning). Most of them use SLIC (Achanta et al. 2012), SAM (Kirillov et al. 2023), and SEEM (Zou et al. 2024) to guide mask segmentation and choose ResNet (He et al. 2016) as the image encoder, which means that mask segmentation and mask feature generation are two completely independent modules. This not only increases the training time but also makes noise easily stack up across different models. At the same time, the lack of language guidance during segmentation will lead to a more random mask granularity. Fortunately, 2D open-vocabulary segmentation models perfectly address this issue, as we can not only extract labels and image embeddings from them but also obtain segmentation with appropriate granularity.
@@
Figure 3: Overview of AFOV, which consists of two stages: Tri-Modal Pre-training (TMP) and Annotation-free training (AFOV-baseline). Both stages leverage masks and mask labels extracted from 2D open-vocabulary segmentation models, while mask features and text features are employed only in TMP. TMP enhances scene understanding through contrastive losses: superpixel-superpoint loss $\mathcal{L}_{I-P}$ and text-superpoint loss $\mathcal{L}_{T-P}$, while our baseline employs pseudo-labels to supervise the 3D network. Additionally, to bridge dataset classes and open vocabularies, we introduce a class dictionary. The Approximate Flat Interaction (AFI) optimizes the results by spatial structural analysis in a broad perception domain.
@@
To summarize, we extract four interrelated, synchronously generated knowledge from CLIP-based open-vocabulary segmentation models: 1) images’ segmentations as masks $M_{\mathcal{I}}$ from image set $\mathcal{I}$; 2) corresponding labels $L_{M}$ for $M_{\mathcal{I}}$; 3) image features corresponding to $M_{\mathcal{I}}$, denoted as $F_{M}$; and 4) text features $F_{T}$. Each of the above knowledge will play an important role in the following sections, as shown in Fig. 3.
Baseline of AFOV
@@
Given a point cloud $\mathcal{P}=\{(p_{n},e_{n})|n=1,\ldots,N\}$, where $p_{n}\in\mathbb{R}^{3}$ represents the 3D coordinates of a point, $e_{n}\in\mathbb{R}^{E}$ denotes the point’s features. ${L}=\{l_{n}|n=1,\ldots,N\}$ are the labels of $\mathcal{P}$ and =$\mathcal{I}=\{i_{k}|k=1,\ldots,K\}$ represents the images captured by a synchronized camera at the same moment. In contrast to supervised methods, our task does not utilize labels ${L}$ during training. We choose to employ a simple way of generating pseudo-labels for point clouds with the assistance of image segmentation: With masks $M_{\mathcal{I}}=\{m_{r}|r=1,\ldots,R\}$ obtained from image set $\mathcal{I}$ as described in the above section, we use the labels ${L}_{M}$ corresponding to $M_{\mathcal{I}}$ as the pseudo-label ${L}_{\text{pixel}}^{\text{pseudo}}$ for pixels in every mask $m_{r}\in M_{\mathcal{I}}$. By leveraging known sensor calibration parameters, we establish a mapping $\Gamma_{\text{camera}\leftarrow\text{LiDAR}}$ to bridge the gap between domains of point clouds and images. Pseudo-labels ${L}_{\mathcal{P}}^{\text{pseudo}}=\{l_{n_{0}}^{\text{pseudo}}|n_{0}=1,\ldots,N_{0}\}$ for point clouds $\mathcal{P}$ are generated through ${L}_{\text{pixel}}^{\text{pseudo}}$ and mapping $\Gamma_{\text{camera}\leftarrow\text{LiDAR}}$. For a 3D backbone $\mathcal{F}_{\theta_{p}}:\mathbb{R}^{N\times(3+L)}\rightarrow\mathbb{R}^{N\times D}$ with the learnable parameter $\theta_{p}$, we train $\theta_{p}$ with pseudo-labels ${L}_{\mathcal{P}}^{\text{pseudo}}$. Given the sparsity of point clouds, it is obvious that $\Gamma_{\text{camera}\leftarrow\text{LiDAR}}$ is not surjective. It is important to note that $\Gamma_{\text{camera}\leftarrow\text{LiDAR}}$ is also not injective, as the projection area of LiDAR is not entirely covered by cameras, resulting in obvious pseudo-label-blank areas in point cloud $\mathcal{P}$. After knowledge distillation, these untrained regions exhibit label confusion, which will be discussed in the next section.
@@
Tri-Modal Contrastive Pre-training (TMP)
In this section, we introduce Tri-Modal contrastive Pre-training (TMP). TMP innovatively integrates textual information into pre-training and removes the 2D backbone through pre-generation of the features, demonstrating excellent performance in both annotation-free training and fine-tuning. The illustration of TMP can refer to Fig. 3.
Synchronous Generation of Knowledge
To the best of our knowledge, most existing 3D pre-training methods (Liu et al. 2021; Sautier et al. 2022; Liu et al. 2023) generate masks and mask features asynchronously for 2D-3D contrastive learning, which causes noise aggregation between different backbones. In TMP, we address this issue by synchronously generating masks and features before pre-training.
Integrating Textual Information in TMP
We aim to incorporate text-3D contrastive learning into pretraining. For different instances of the same category (e.g., Car A and Car B), they share the same textual features but have different image features. Theoretically, 2D-3D contrastive learning provides rich features from the instance level (needed for pretraining); text-3D contrastive learning, on the other hand, offers direct semantic embeddings for the 3D backbone from the semantic level (helpful for annotation-free segmentation). So the design intention of TMP is: Can integrating textual information into pre-training accomplish multiple tasks (pretraining and unannotated segmentation)? Experiment removing text-3D contrastive learning is conducted in Appendix B.
Figure 4: Illustrating two examples of potential ”self-conflicts” based on SAM segmentation.
”Self-Conflict”
Regions with the same semantics may be divided into multiple parts (in 2D or 3D). During contrastive learning, different parts act as ”negative samples” to each other, causing features of parts with the same semantics to be pushed apart in the feature space. This phenomenon is called ”self-conflict”. The usage of superpixels and superpoints for segmentation in SLidR (Sautier et al. 2022) is inspiring, which partly alleviates the issue of ”self-conflict” caused by point-level point set partitioning (e.g. $k$-NN), such as ”negative sample 1” in Fig. 4. In Appendix B, we conduct an ablation study on the use of superpixels and superpoints, and the results demonstrate their effectiveness.
However, the ”self-conflict” caused by the mask granularity randomness of some VLMs has not been properly addressed, such as ”negative sample 2” in Fig. 4. For instance, SAM (Kirillov et al. 2023) might separate the car-door and car-window because SAM is unaware of the appropriate segmentation granularity (both should belong to the same class: ”car”). This issue arises due to the lack of textual assistance. When guided by the text, a 2D open-vocabulary segmentation model would treat the entire car as a whole, thereby avoiding ”self-conflict”.
Superpixel-Superpoint Generation